# Tactile SoftHand-A: 3D-Printed, Tactile, Highly-underactuated, Anthropomorphic Robot Hand with an Antagonistic Tendon Mechanism

### v1

@@ 2.2 Vision-based Tactile Sensing
Compared to piezoelectric and resistive sensors, vision-based tactile sensors (VBTS) are increasingly favored by robotics experts due to their superior spatial resolution, which provides richer physical information for robotic interaction tasks, such as texture features and object pose. Researchers have developed many VBTS, such as MIT’s Gelsight (Yuan et al. (2017)), META’s DIGIT (Lambeta et al. (2020)), BRL’s TacTip (Ward-Cherrier et al. (2018), Lepora (2021)) and BioTacTip (Li et al. (2024)). However, the manufacturing processes for these sensors are complex and labor-intensive, typically exceeding five hours. For instance, creating the DIGIT involves using specialized molds for silicone casting to form its soft skin. Also, the use of transparent windows in all VBTSs mentioned above requires laser equipment for cutting and hand-placement/gluing the window. In manufacturing the TacTip and BioTacTip(Li et al. (2024)), transparent gel is injected into the skin and window cavities, and specialized vacuum equipment is necessary to remove any bubbles that may form during injection and curing. Furthermore, when integration with a robotic fingertip is needed, researchers must create new molds, and produce connection parts to the phalanges. Additionally, the manual production methods of VBTS, such as variations in silicone pouring and injection, lead to physical performance discrepancies among sensors, such as stiffness and sensor skin shape, which pose challenges in generalizing models trained on individual sensors in practical applications.
@@
Figure 3: Antagonistic tendon comparison for P-type, D-type and A-type fingers. (A) Side view of three type fingers. (B) Schematic model of all fingers showing the main variables of interest (described in TABLE I).
@@ 3 Development of Robot Fingers Featuring Active Antagonism
Based on different tendon configurations, we have developed three types of fingers, each exhibiting distinct flexion gestures, termed the A-type finger, D-type finger and P-type finger, respectively (for antagonistic, distal and proximal). The A-type finger, which features an antagonistic tendon running through the entire finger, exhibits the fastest extension. The D-type finger features active control of the distal interphalangeal (DIP) joint, achieved through the coordination of a driving tendon and an antagonistic tendon and a passive element on the DIP joint. Similarly, the P-type finger employs a comparable mechanism to actively control the proximal interphalangeal (PIP) joint. Their design details are as follows.
3.1 Design of A-type finger
@@
In contrast to the D-type and P-type fingers, the antagonistic tendon of the A-type finger terminates at the fingertip (Figure 4(A)). This means that the A-type finger does not require an elastic band or other passive element to assist with its extension, resulting in a simpler structure. Consequently, during finger flexion, the absence of resistance from an elastic band leads to greater fingertip force and faster reactions compared to the other two types of fingers. However, despite these advantages, the A-type finger lacks active control capabilities for the DIP or PIP joints, unlike the other two types, the effect of which will be assessed experimentally.
3.2 Design of D-type finger
@@
As illustrated in Figure 2(A), the D-type finger primarily consists of phalanges, phalangeal coverings, tendons, and a bearing mechanism. Its total length, width, and height are 115 mm, 20 mm, and 23 mm, respectively. The phalanges are composed of a fingertip, two middle phalanges, and a base phalanx, interconnected through gear engagement to form three rotary joints: the Distal Interphalangeal (DIP), Proximal Interphalangeal (PIP), and Metacarpophalangeal (MCP) joints. A bearing mechanism with multiple U-type grooves is also affixed to the four phalanges. The driving tendon (Figure 2(A), red cable) and antagonistic tendon (Figure 2(A), blue cable) run through the entire finger via the U-type grooves on the bearings. For instance, bearings arranged in an equilateral triangle on the right side of the middle phalanx facilitate flexion movement by providing driving force to the driving tendon. Conversely, bearings in an inverted triangle arrangement on the left side support extension movement. Notably, the antagonistic tendon of the D-type finger terminates at the first middle phalanx. Therefore, the force required for resetting the DIP joint is provided by an elastic band positioned between the fingertip and the first middle phalanx, coupled to the active reset of the other finger joints.
3.3 Design of P-type finger
The primary structure of the P-type finger is identical to that of the D-type finger. The only difference lies in the configuration of the antagonistic tendon (see Figure 4). Unlike the D-type finger, the antagonistic tendon of the P-type finger terminates at the second middle phalanx. Consequently, it is necessary to incorporate passive components at the DIP and PIP joints to facilitate the repositioning of the corresponding phalanges, where we use elastic bands similar to the D-type finger.
3.4 Dynamics Analysis of D-type Finger
This section primarily analyzes how to achieve control of the DIP joints of the fingers at any position through coordinated control of the driving tendons and antagonistic tendons. Several methods exist for modeling the dynamics of tendons, including particle models, continuum models, and rigid body models. However, for the Tactile SoftHand-A, which features ten tendons connected by differential mechanisms and passive components, such models become overly complex and difficult to integrate with control systems. Therefore, we adopt an analytic model in which the force distribution and the schematic of the finger are shown in Figures 2(D) and 4(B) respectively. The meaning of all notation for finger dynamics analysis are introduced in TABLE I. The meanings of all notations used for finger dynamics analysis are presented in TABLE I. There are two periods that describe the movement of the D-type finger: one without active control of the DIP joint from time $t_{0}$ to $t_{1}$, and another with active control of the DIP joint from $t_{1}$ to $t_{2}$.
The finger rotation angle $\theta(t)$ is given by the sum of DIP rotation angle $\theta_{d}(t)$, PIP rotation angle $\theta_{p}(t)$ and MCP rotation angle $\theta_{m}(t)$:
$\theta(t)=\theta_{m}(t)+\theta_{p}(t)+\theta_{d}(t).$ (1)
The change of driving tendon and antagonistic tendon can be represented by $\Delta{l_{d}}(t)$ and $\Delta{l_{a}}(t)$, where
$\Delta{l_{d}}(t)=\left|{l_{d}}(t)+{l_{d}}(t_{0})\right|,$ (2)
$\Delta{l_{a}}(t)=\left|{l_{a}}(t)+{l_{a}}(t_{0})\right|.$ (3)
@@ 3.4 Dynamics Analysis of D-type Finger
First, we consider a period from $t_{0}$ to $t_{1}$, in which the active antagonism synchronizes with the driving tendon. Therefore, there is no resistance acted on MCP and PIP joint:
$\tau_{m\_a}(t_{1})=\tau_{p\_a}(t_{1})=0,$ (4)
$\Delta{l_{d}}(t_{1})$ $=\Delta{l_{a}}(t_{1})=\theta(t_{1})\cdot l$ (5)
$=\theta_{m}(t_{1})\cdot l_{3}+\theta_{p}(t_{1})\cdot l_{2}+\theta_{d}(t_{1})\cdot l_{1}.$
Now, in a second period from $t_{1}$ to $t_{2}$, we use active antagonism to control the DIP joint. We keep the active antagonism at the same length and pull the driving tendon:
$\Delta{l_{a}}(t_{2})=\Delta{l_{a}}(t_{1}),\ \ \tau_{m\_a}(t_{2})=\tau_{p\_a}(t_{2})=\infty.$ (6)
Therefore, the difference in tendon lengths between the driving and antagonistic tendon is
$\Delta{l_{d}}(t_{2})-\Delta{l_{a}}(t_{2})$ $=\Delta{l_{d}}(t_{2})-\Delta{l_{a}}(t_{1})$ (7)
$=(\theta_{d}(t_{2})-\theta_{d}(t_{1}))\cdot l_{1},$
@@
Therefore, we can actively control the DIP joint by controlling the displacement of driving tendon and active antagonistic tendon.
@@
3.5 Comparisons of Finger Workspace in Simulation
After demonstrating that A-type, D-type and P-type fingers can execute diverse gestures through the coordinated control of driving and antagonistic tendons, an important comparison is their workspaces. To this end, we utilized the MATLAB Robotics Toolbox to construct models for these finger types, employing the Denavit-Hartenberg (DH) matrix to evaluate their workspaces. For example, the D-type finger, requires an independent control input for active management of its DIP joint due to its capability for autonomous movement; meanwhile, the proximal PIP and MCP joints are interconnected, necessitating a separate control input to adjust the angles of both PIP and MCP joints. By applying the Monte Carlo method to generate varied inputs, we employ forward kinematics to delineate the fingertip’s workspace (Figure 4).
Using the above method, we plotted the fingertip vertex positions for various gestures in the simulator (Figure 4). Unlike the D-type and P-type fingers, A-type fingers lack active control over the DIP or PIP joints, resulting in a workspace confined to a fixed arc due to the coordinated action of driving and active antagonistic tendons (Figure 4(c)). Notably, the D-type finger demonstrates a more uniform workspace distribution, particularly in extended gestures, leading to the selection of D-type configured fingers for testing and developing the Tactile SoftHand-A, as discussed in the later sections of this paper.
It is worth noting that the above simulation was completed without touching any objects or obstacles; because the fingers have adaptive capabilities, simulating workspace changes in contact with objects or obstacles are not discussed in this section, but instead are evaluated experimentally later.
4 Development of fully-integrated tactile fingertips
The fingers are intended to have optical tactile sensing in the distal phalanges. Here we introduce a novel design where the entire sensing structure of the fingertip is 3D-printed as a single component integrated into the physical structure of the finger, requiring no manual fabrication except for the insertion of the camera module. First, we detail the main components and design considerations of the fully-integrated fingertips, then their fabrication utilizing advances in multi-material 3D-printing technology.
4.1 Integrated tactile fingertip design
@@
The perception and actuation modules are also mounted on the phalangeal skeleton, with bearings and the tendon mechanism secured by bolts on both sides. The camera lens and camera board are affixed to the top of the fingertip. The skin surface, located at the fingertip’s pad, converts external contacts into internal deformation, thereby moving markers attached on pins to its inner surface based on the TacTip design (Ward-Cherrier et al. (2018)). These are then captured by the camera to form tactile images. Additionally, the black outer skin surface and window cover isolate external light interference, to aid producing high-quality tactile images. The soft, transparent clear fill is designed to increase the surface’s deformation range while providing internal support and resetting its deformation. The rigid, transparent clear window serves to support the clear fill and refract light from the LED strip.
@@ 4.2 Fabrication of the tactile fingertip
The perception and actuation module components of the tactile fingertip can be purchased inexpensively, while the remaining components are all produced using multimaterial 3D-printing (Stratasys J826). The entire assembly of the contact module’s components, including the soft black opaque surface (soft agilus material), hard white markers (hard verowhite material), soft transparent filling (soft agilus clear mixed with support material), hard transparent window (hard veroclear material), hard black opaque window cover, and the hard white skeleton (hard verowhite material), are integrally formed as a single component without any other post-fabrication processes required.
@@
This capability is made possible by an innovation we have introduced in mixing support material with Agilus material to print a clear fill that, while maintaining its transparency, can have its hardness adjusted by altering the proportion of support material. Previously, the hardness of the pure Agilus material available with the printer was found to be excessively high for use as a soft filler. Usually, the support material is required for the layering of complex structures and is cleaned post-printing. Fortunately, one of the support materials happens to be optically clear and we found it is compatible with mixing with other materials, which lends it to adjusting the filling material’s softness.
@@ 4.3 Tactile Localization Model
The camera images through the perception model are then processed with a corresponding tactile model (Figure 6). Here we are primarily concerned with predicting and using the contact region and its centre point, which we will integrate into the robot hand control. The model details are described below.

### v2

@@ 2.2 Tactile Sensing in Robot Hands
However, the manufacturing processes for these vision-based tactile sensors are complex and labor-intensive, typically exceeding many hours (Zhang et al. (2022)). Many such sensors derived from the GelSight, such as the 9DTact (Lin et al. (2023)), require specific molds and the casting of silicone to create soft skins. Also, the use of transparent windows in all VBTSs mentioned above requires laser equipment for cutting and hand-placement/gluing the window. In manufacturing the TacTip, transparent gel is injected into the skin and window cavities, and specialized vacuum equipment is necessary to remove any bubbles that may form during injection and curing (Ward-Cherrier et al. (2018)). Furthermore, when integration within robot hands is needed, researchers must create new molds and produce connection parts to the phalanges or palm (Zhang et al. (2024b, c)). These manual manufacturing processes significantly restrict the form-factor of vision-based tactile sensors, as well as the variety and distribution of inner structures such as markers. Furthermore, the manufacturing variation inherent in manual production can impact the generalization performance of perception systems using the tactile images to predict contact features, as the perception models can vary even across fingertips on the same robot hand (Ford et al. (2025)).
@@
Figure 3: Antagonist tendon comparison for P-type (proximal-type), D-type (distal-type) and A-type (antagonstic-type) fingers. (A) Side views. (B) Kinematically-equivalent model of D-type finger showing the main variables of interest (described in inset Table).
@@ 3 Development of Robot Fingers Featuring Active Antagonism
Based on different tendon configurations, we have developed three types of fingers, each exhibiting distinct flexion gestures, termed the A-type finger, D-type finger and P-type finger for fully antagonistic-type (A), distal-type (D) and proximal-type (P), respectively (see Figures 4(A), 4 and 8). The A-type finger, which features an antagonist tendon running through the entire finger, exhibits the fastest extension. The D-type finger features active control of the distal interphalangeal (DIP) joint, achieved through the coordination of agonist and antagonist tendons and a passive element on the DIP joint. Similarly, the P-type finger employs a comparable mechanism to actively control the proximal interphalangeal (PIP) joint. Their design details are as follows.
3.1 Design of A-type (antagonistic-type) finger
As shown in Figure 2(D), the bearings used for winding the agonist tendon are arranged in an upright triangular configuration, while the bearings for winding the antagonist tendon are arranged in an inverted triangular configuration. This arrangement is designed to generate a greater downward force on the tendons in the middle pulley, thereby applying sufficient torque to the joints to drive joint rotation or reset the joint. In this study, the pulley arrangement is the same for all types of fingers, with the height of the middle pulleys on each phalanx kept consistent. Due to friction and tendon deformation, the MCP joint typically experiences the highest torque first, resulting in its rotation, which provides the finger with a larger workspace and greater adaptability (this phenomenon is demonstrated in the multimedia video file). To achieve specific finger gestures (e.g. the synchronous rotation of all joints in Megaro et al. (2017)), it is necessary to adjust the height of the middle pulleys on each phalanx. The arrangement of the pulley system for the antagonist tendons follows the same principle and similar considerations apply.
Similarly to the design of the BPI-SoftHand (Li et al. (2022)), a connecting rod at each joint links the adjacent phalanges to maintain correct gear engagement and prevent dislocation (differing from the original Pisa/IIT-SoftHand design of Catalano et al. (2014)).
@@ 3.1 Design of A-type (antagonistic-type) finger
In contrast to the D-type and P-type fingers, the antagonist tendon of the A-type finger terminates at the fingertip (Figure 4(A)). This means that the A-type finger does not require an elastic band or other passive element to assist its extension, resulting in a simpler structure and full active antagonism. Consequently, during finger flexion, the absence of resistance from an elastic band gives a greater fingertip force and faster reaction compared to the other two types of fingers. However, despite these advantages, the A-type finger lacks active control capabilities for the DIP or PIP joints, unlike the other two types, the effect of which is assessed experimentally below.
3.2 Design of D-type (distal-type) finger
@@
As illustrated in Figure 2(A), the D-type finger primarily consists of phalanges, phalangeal coverings, tendons, and a bearing mechanism. The phalanges are composed of a fingertip, two middle phalanges, and a base phalanx, interconnected through gear engagement to form three rotary joints: the Distal Interphalangeal (DIP), Proximal Interphalangeal (PIP), and Metacarpophalangeal (MCP) joints. A bearing mechanism with multiple U-type grooves is also affixed to the four phalanges. The agonist tendon (Figure 2(A), red cable) and antagonist tendon (Figure 2(A), blue cable) run through the entire finger via the U-type grooves on the bearings. For instance, bearings arranged in an equilateral triangle on the right side of the middle phalanx facilitate flexion movement by providing agonist force to the agonist tendon. Conversely, bearings in an inverted triangle arrangement on the left side support extension movement. Notably, the antagonist tendon of the D-type finger terminates at the first middle phalanx. Therefore, the force required for resetting the DIP joint is provided by an elastic band positioned between the fingertip and the middle phalanx, coupled to the active reset of the other finger joints.
3.3 Design of P-type (proximal-type) finger
The primary structure of the P-type finger is identical to that of the D-type finger. The only difference lies in the configuration of the antagonist tendon (see Figure 4). Unlike the D-type finger, the antagonist tendon of the P-type finger terminates at the second middle phalanx. Consequently, it is necessary to incorporate passive components at the DIP and PIP joints to facilitate the repositioning of the corresponding phalanges, where we use elastic bands similar to the D-type finger. It is worth noting that the position of the elastic bands significantly affects the amount of resistance they generate at the joints, which influences the gesture of the finger. For the D-type finger, the DIP joint is fully decoupled and thus remains unaffected by the elastic bands. However, for the P-type finger, the PIP and DIP joints are still connected via elastic bands, meaning that the relative rotation between these joints is influenced by the stiffness of the elastic bands. The farther the elastic bands are positioned from the joint’s rotational center, the greater their deformation during rotation, which in turn increases the resistance applied to the joint. Therefore, by adjusting the position of the elastic bands on DIP or PIP joint independently, the joint resistance can be set for precise regulation of the finger’s flexion gesture.
3.4 Kinematic Analysis of D-type Finger
The D-type finger design features active control of the DIP joints of the fingers at any position through coordinated control of the agonist and antagonist tendons. To examine this feature, we adopt an analytic model in which the force distribution and the schematic of the finger are shown in Figures 2(D) and 4(B) respectively, with symbol notation introduced in Table I. We consider two time periods for the movement of the D-type finger: the first without active control of the DIP joint from time $t_{0}$ to $t_{1}$, and the second with active control of the DIP joint from $t_{1}$ to $t_{2}$.
The finger rotation angle $\theta(t)$ is given by the sum of DIP rotation angle $\theta_{\rm d}(t)$, PIP rotation angle $\theta_{\rm p}(t)$, and MCP rotation angle $\theta_{\rm m}(t)$:
$\theta(t)=\theta_{\rm m}(t)+\theta_{\rm p}(t)+\theta_{\rm d}(t).$ (1)
Accordingly, the length change of the agonist and antagonist tendons is represented by $\Delta{l_{\rm d}}(t)$ and $\Delta{l_{\rm a}}(t)$, where
$\Delta{l_{\rm d}}(t)=\left|{l_{\rm d}}(t)-{l_{\rm d}}(t_{0})\right|,$ (2)
$\Delta{l_{\rm a}}(t)=\left|{l_{\rm a}}(t)-{l_{\rm a}}(t_{0})\right|.$ (3)
@@ 3.4 Kinematic Analysis of D-type Finger
First, let us consider a period from $t_{0}$ to $t_{1}$, in which the antagonism synchronizes with the agonist tendon. Therefore, there is no resistance acting on the MCP and PIP joints:
$\tau_{\rm a}^{\rm mcp}(t_{1})=\tau_{\rm a}^{\rm pip}(t_{1})=0,$ (4)
$\Delta{l_{\rm d}}(t_{1})-\Delta{l_{\rm a}}(t_{1})=\theta_{\rm d}(t_{1})r.$ (5)
Now, in a second period from $t_{1}$ to $t_{2}$, we use active antagonism to control the DIP joint. We keep the antagonist tendon at the same length and pull the agonist tendon. This means that the torque acting on the MCP and PIP joints by the agonist tendon is equal to the torque acting on the MCP and PIP joints by the antagonist tendon, and is less than the maximum stall torque of the antagonist motor, $\tau_{\rm a}^{\rm stall}$. Thus, the MCP and PIP joints will remain in their current state:
$\tau_{\rm d}^{\rm mcp}(t_{1})+\tau_{\rm d}^{\rm pip}(t_{1})=\tau_{\rm a}^{\rm mcp}(t_{1})+\tau_{\rm a}^{\rm pip}(t_{1})\leq\tau_{\rm a}^{\rm stall},$ (6)
$\theta_{\rm m}(t_{1})\!=\!\theta_{\rm m}(t_{2}),\,\theta_{\rm p}(t_{1})\!=\!\theta_{\rm p}(t_{2}),\,\Delta{l_{\rm a}}(t_{2})\!=\!\Delta{l_{\rm a}}(t_{1}).$ (7)
Hence, the difference in length of the agonist tendon between $t_{1}$ and $t_{2}$ is
$\Delta{l_{\rm d}}(t_{2})-\Delta{l_{\rm d}}(t_{1})$ $=(\theta_{\rm d}(t_{2})-\theta_{\rm d}(t_{1}))r,$ (8)
@@
Therefore, we can actively control the DIP joint by controlling the displacement of the agonist and antagonist tendons.
@@
3.5 Simulation of Finger Workspace
After demonstrating that the A-type, D-type, and P-type fingers can execute diverse gestures through the coordinated control of the agonist and antagonist tendons, an important comparison is their workspaces. To this end, we utilized the MATLAB Robotics Toolbox to construct models for these finger types, employing the Denavit-Hartenberg (DH) matrix. This serves as a preliminary step to validate the overall design concept and predict the workspace of different finger-type configurations. This enables us to systematically compare the workspaces of different finger types under the same conditions, which would otherwise be highly time-consuming and challenging to implement using physical prototypes for each design iteration.
For example, the D-type finger requires an independent control input for active management of its DIP joint due to its capability for autonomous movement; meanwhile, the proximal PIP and MCP joints are interconnected, necessitating a separate control input to adjust the angles of both PIP and MCP joints. By applying the Monte Carlo method to generate varied inputs, we employ forward kinematics to delineate the fingertip’s workspace.
Using the above method, we plotted the fingertip vertex positions for various gestures in the simulator (Figure 4). Unlike the D-type and P-type fingers, the A-type fingers lack active control over the DIP or PIP joints, resulting in a workspace confined to a fixed arc due to the coordinated action of the agonist and antagonist tendons (Figure 4(c)). Notably, the D-type finger demonstrates a more uniform workspace distribution, particularly in extended gestures, leading to the selection of D-type configured fingers for testing and developing the Tactile SoftHand-A, as discussed in the later sections of this paper.
It is worth noting that the above simulation was completed without touching any objects or obstacles. Because the fingers have adaptive capabilities, simulating workspace changes in contact with objects or obstacles is not discussed in this section, but is evaluated experimentally later instead.
4 Development of Integrated Tactile Fingertips
The fingers are intended to have vision-based tactile sensing in the distal phalanges. Here we introduce a novel design where the entire sensing structure of the fingertip is 3D-printed as a single component integrated into the physical structure of the finger, requiring no manual fabrication except for the insertion of the camera module. First, we detail the main components and design considerations of the fully-integrated fingertips, then their fabrication utilizing advances in multi-material 3D-printing technology.
4.1 Overall Design
@@
The perception and actuation modules are also mounted on the phalangeal skeleton, with bearings and the tendon mechanism secured by bolts on both sides. The camera lens and camera board are affixed to the top of the fingertip. The skin surface, located at the fingertip’s pad, converts external contacts into internal deformation, thereby moving markers attached on pins to its inner surface based on the TacTip design (Ward-Cherrier et al. (2018); Lepora (2021)). These are then captured by the camera to form tactile images. Additionally, the black outer skin surface and window cover isolate external light interference, to aid in producing high-quality tactile images. The soft, transparent clear fill is designed to increase the surface’s deformation range while providing internal support and resetting its deformation. The rigid, transparent clear window serves to support the clear fill and refract light from the LED strip.
@@ 4.2 Fabrication of the Tactile Fingertip
The perception and actuation module components of the tactile fingertip can be purchased inexpensively, and the remaining components produced using multimaterial 3D-printing (Stratasys J826 Prime 3D-printer). The entire assembly of the contact module’s components, including the flexible black opaque surface and pins (soft Agilus30 material), hard white markers (hard VeroWhite material), soft transparent filling (soft Agilus30 clear mixed with support material), hard transparent window (hard VeroClear material), hard black opaque window cover, and the hard white skeleton (hard VeroWhite material), is integrally formed as a single component without any other post-fabrication processes required.
@@
This capability is made possible by an innovation we have introduced in mixing support material with Agilus material to print a clear fill that, while maintaining its transparency, can have its hardness adjusted by altering the proportion of support material. Previously, the hardness of the pure Agilus material available with the printer was found to be excessively high for use as a soft filler. Usually, the support material is required for the layering of complex structures and is cleaned post-printing. We noticed that one of the support materials for complex pieces happens to be optically clear, and through experimenting, found it is compatible with mixing with other materials. Hence, we use the support material to adjust the filling material’s softness to a level that would otherwise not be possible using commercial 3D-printing materials.
Other possibilities to extend these techniques would be to use solid-liquid hybrid 3D printing, such as proposed by MacCurdy et al. (2016). However, those methods use liquid fillers and require additional support materials to prevent interaction between uncured liquid and the deposited photopolymer, which is somewhat more complicated than our approach. With advances in 3D-printing technology, we expect better techniques to become possible for fabricating diverse optical tactile sensors.
@@ 4.3 Tactile Localization Model
The camera images captured with the perception module are then processed with a tactile model (Figure 6). Here we are primarily concerned with predicting and using the contact region and its center point, which we will integrate into the robot hand control. The model details are described below.